3.2.29 Scores of an IQ test have a bell-shaped distribution with a mean of 100 and a standard deviation of 14. Use the empirical rule to determine the following. (a) What percentage of people has an IQ score between 72 and 128? (b) What percentage of people has an IQ score less than 72 or greater than 128? (c) What percentage of people has an IQ score greater than 142? (a)% (Type an integer or a decimal.)
Answer #1

a)as 72 and 128 are 2 std deviation from mean ; therefore 95%

b) percentage of people =(100-95)% =5%

c) as 142 is 3 std deviation from mean; hence % of people =(100-99.7)/2 =0.15 %